(7 pts.) Let f(x, y, z) = y and let R be the region {(x, y, z) |2 < x < 4,0 Sy < 3,15 zse}. 2 Evaluate | $180,0,.2) av. R
Answer #1

Solution: - Guiven 4,2) av Solution: - Gown SSS Fem 19, 2) av where flavy,z) = xy2 i and R. be the region I cany, z)/2 <a <4,خاله و تور در dz dy /\/ -7 اور ) == له رقم - ته) - ه - ه . 1=In ] بده = = 54 2 we - In] رہ- () 1 5 = 59 = 0- ا5 =